Define f(x) = sinx on [0, 2pi]. Find two increasing functions h and g for which f = h — g on 

[0, 2pi]. 


1
Expert's answer
2021-02-24T07:41:17-0500

"\displaystyle

f = \sin{x} = \cos\left(\frac{\pi}{2} - x\right) \\

f = h \mapsto g \\

h(x) = \frac{\pi}{2} - x,\,\,g(x) = \cos{x}"